Circular No. 7559

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V445 PUPPIS N. N. Samus, Sternberg Astronomical Institute, Moscow, informs us

that the designation V445 Pup has been given to the possible nova reported on IAUC 7552 and 7556.

N. M. Ashok and D. P. K. Banerjee, Physical Research Laboratory, Ahmedabad, report 1.1–2.3-μm spectroscopy of V445 Pup, with the Mt. Abu 1.2-m telescope (+ PRL Near-Infrared Imaging Spectrometer + HgCdTe NICMOS3 detector array) on Jan. 2.875 UT: “The J- and Hband (1.08–1.78 μm) spectra are dominated by emission features arising from C I, C III, and C IV (the prominent ones being at 1.167, 1.177, 1.259, 1.371, and 1.691 μm). A fairly strong unidentified emission feature is seen at 1.132 μm. The Brackett-series lines of H I in the K and H bands and the Paβ line are not present. However, an emission feature at 1.095 μm, coinciding with the Paγ line, is likely due to He II. A weak emission line with a P-Cyg profile at 2.062 μm is attributable to He I. A discernible rise in the continuum is seen redward from 1.60 μm, indicating the presence of a cooler component.”

T. Kato, Kyoto University, reports that a 14-min Tri-X exposure taken by Y. Nakamura (Suzuka, Mie, Japan, 200-mm f/4 camera lens) on 2000 Sept. 26.78 UT shows no object to mag 12.0 at the location of V445 Pup.

Photometry by A. C. Gilmore with the 0.6-m f/16 reflector at the University of Canterbury’s Mt. John Observatory, Lake Tekapo, New Zealand (reference star Cousins E352; uncertainties ∼ 0.01 mag): Jan. 6.58 UT, V = 9.05, U − B = −0.44, B − V = +0.62, V − R = +0.40, V − I = +1.04.

COMET C/1999 T1 (McNAUGHT-HARTLEY) N. Biver, D. Bockelée-Morvan, and J. Crovisier, Observatoire de Paris-

Meudon; D. C. Lis, California Institute of Technology; and H. Weaver, Johns Hopkins University, report: “The CO J(3–2) line at 345.8 GHz has been detected on Jan. 5.7 UT at the Caltech Submillimeter Observatory (CSO) with a line area of 0.17 ± 0.03 K km/s in main-beam brightness temperature. During Jan. 5–7 at CSO, we also detected the HCN J(3–2) line (0.19 ±0.01 K km/s) and CH_{3}OH lines at 307.2 (0.23 ± 0.02), 304.2 (0.16 ± 0.02), and 241.79 GHz (0.17 ± 0.03 K km/s). The average production rates relative to water, using the water-production rate from Schleicher et al. (cf. IAUC 7558), are: CO, 40 percent; CH_{3}OH, 5 percent; HCN, 0.11 percent. This is until now the highest mixing ratio of CO observed in a comet that is relatively close to the sun.”

2001 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green